An object is propelled vertically upward with an initial velocity of 20 meters per second. The distance s (in meters) of the object from the ground after t seconds is s = -4.9t^2 + 20t.
(a) When will the object be 15 meters above the ground?
(b) When will it strike the ground?
(c) Will the object reach a height of 100 meters?

Answer by stanbon(57323)

An object is propelled vertically upward with an initial velocity of 20 meters per second. The distance s (in meters) of the object from the ground after t seconds is s(t) = -4.9t^2 + 20t.
---------
The equation tells you how far above the ground the object after "t" seconds.
I'll just point out the equations you should solve:
--------------------
(a) When will the object be 15 meters above the ground?
Solve -4.9t^2 + 20t = 15
You should get two solutions: one for the time as it rises and one for
the time it is decending..
--------------------------------------
(b) When will it strike the ground?
Solve -4.9t^2 + 20t = 0
You should get twon solutions. One of the solutions is t=0.
------------------------------------
(c) Will the object reach a height of 100 meters?
Solve -4.9t^2 + 20t = 100
Two solutions.
=================
Cheers,
Stan H.